OpenStudy (anonymous):

If f(x) = x^3/2, what is (f.f)(-4)?

OpenStudy (anonymous):

is (f.f), f times f or f of f?

OpenStudy (anonymous):

f of f

OpenStudy (anonymous):

(f of f)(x)= (x^3/2)^3/2 sub x=-4 (f of f)(-4)=-16384

OpenStudy (anonymous):

if f(x)= { (-7,1),(-5,8} and g= { (-6,4), (-5,3)} how to find (f+g)?

OpenStudy (anonymous):

Actually there is more numbers for f(x) and g(x)

OpenStudy (anonymous):

f(x)= { (-7,1),(-5,8),(3,11),(5,-1) and g(x) {(-6,4),(-5,3),(-1,7),(5,5)}

OpenStudy (anonymous):

I need example to slove this question. Do u know how?

OpenStudy (anonymous):

just add the y values for the matching x values and keep the rest (assuming f and g are sets) so (f+g)={(-7,1),(-6,4),(-5,11),(-1,7),(3,11),(5,4)}

OpenStudy (anonymous):

sorry I don't get it.

OpenStudy (anonymous):

so f contains (-5,8) and g also contains (-5,3), since x=-5 appears in both, f+g for that point equals (-5,8+3)=(-5,11)

OpenStudy (anonymous):

thx i got it

OpenStudy (anonymous):

you either keep all the rest of the points that don't have the same x values or don't keep them at all depending on the behaviour of the function which is not defined in the question

OpenStudy (anonymous):

i gonna solve other question and i will add another question. Thank for your help

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

f(x)= /3x/-10 and g(x)=2^3, what is (g.f)(-2)?

OpenStudy (anonymous):

i have confuse about /3x/ ?

OpenStudy (anonymous):

|x| means what ever value in |x| is positive example |-2|=2 i think you typed g(x) wrong since there is no variable

OpenStudy (anonymous):

| | means absolute value

OpenStudy (anonymous):

|3x| means absolute value of 3x

OpenStudy (anonymous):

g(x)=2x^3

OpenStudy (anonymous):

f(-2)=|3(-2)|-10=-4 (g of f)(-4)=2f(-2)^3=-128

OpenStudy (anonymous):

Thx